Fyodor Corporation has a Parts Division that does work for other Divisions in the company as well as for outside customers. The company's Machine Division has asked the Parts Division to provide it with 9600 special parts each year. The special parts would require $39 per unit in variable production costs.

The Machine Division has a bid from an outside supplier for the special parts at $55.50 per unit. In order to have time and space to produce the special part, the Parts Division would have to cut back production of another part-the QR4 that it presently is producing. The QR4 sells for $76 per unit, and requires $38 per unit in variable production costs. Packaging and shipping costs of the QR4 are $2 per unit. Packaging and shipping costs for the new special part would be only $0.50 per unit. The Parts Division is now producing and selling 48,000 units of the QR4 each year. Production and sales of the QR4 would drop by 5% if the new special part is produced for the Machine Division.

Required:

Please calculate the following:

Fyodor Parts Division Selling transfer price (rounded to 2 decimal places)

$

Fyodor Machine Division Buying Transfer price (rounded to 2 decimal places)

$

Should the transfer take place (Yes or No)
Answer #1
Given,
Variable cost to produce special parts = 39
Parts Division selling price = variable cost + specific fixed cost + oppurtunity cost foregone to produce the product
Since, the Parts division has to reduce the production of QR4 to produce the special parts (Parts Division operating at full capacity) oppurtunity cost is
Selling price of QR4 = 76
Variable cost of QR4 = 38
Packing cost of QR4 = 2
Contribution of QR4 = 76-38-2 = 36
Reduction in production of QR4 on account of producing special parts = 48,000 * 5% = 2,400
Oppurtunity cost = 2,400 *36 = 86,400
Parts Division selling price = variable cost + specific fixed cost + oppurtunity cost foregone to produce the product
                        = [9,600*(39+0.50)]+0+86400
                        = 4,65,600
Selling price / unit = 4,65,600/9,600 =48.5
Machine Division buying price is lower of
(i) External market price 55.5
(ii) Market price of substitutes                -  
(iii) Buying departments ability to pay
(Buying dept selling price - variable cost of buying dept. - specific fixed cost)
               -  
So, machine division buying price = 55.5
Since, Parts Division selling price is less than Machine Division buying price, transfer should take place. It benefits the company.